OpenStudy (anonymous):

Determine whether the function g(x)={(−2x + 10) x≤4 (x^2-3) x>4 is continuous at 4?

OpenStudy (anonymous):

using limit in calculus \[\lim _{x->4^{-}}-2x +10= -2(4)+10= 2\] \[\lim _{x->4^{+}} x ^{2} -3= 16-3=11\] Two limits are different, so this is not continuous

OpenStudy (anonymous):

thank you!